# TinkerXR: In-Situ, Reality-Aware CAD and 3D Printing Interface for Novices

### v3

@@ Abstract.
Despite the growing accessibility of augmented reality (AR) for visualization, existing computer-aided design (CAD) systems remain confined to traditional screens or require complex setups or predefined parameters, limiting immersion and accessibility for novices. We present TinkerXR, an open-sourced interface enabling in-situ design and fabrication through Constructive Solid Geometry (CSG) modeling. TinkerXR operates solely with a headset and 3D printer, allowing users to design directly in and for their physical environments. By leveraging spatial awareness, depth occlusion, recognition of physical constraints, reference objects, and intuitive hand movement controls, TinkerXR enhances realism, precision, and ease of use. Its AR-based workflow integrates design and 3D printing with a drag-and-drop interface for a 3D printer’s virtual twin. A user study comparing TinkerXR with TinkerCAD demonstrates higher accessibility, engagement, and ease of use for novices. By bridging the gap between digital creation and physical output, TinkerXR transforms everyday spaces into accessible and expressive creative studios.
@@
journalyear: 2025copyright: rightsretainedconference: arXiv; preprint; 2025booktitle: arXiv
Figure 1. TinkerXR is an in-situ CAD and 3D printing AR interface. The figure illustrates key features such as spatial awareness, depth occlusion, and the use of physical objects and constraints for design. With Constructive Solid Geometry modeling, users can convert surfaces into interactive design canvases, and after designing, interact with 3D printers’ digital twins to prepare and initiate fabrication. The example designs include a pen holder, phone stand, cable organizer, plant pot holder, bottle mount, and key hook.
[…]
Five images labeled (a), (b), (c), (d), and (e) show the user interacting with the system. In pictures (a), (b), and (c), the user is seen wearing a headset. In picture a, the user is in a room with a desk. The user is standing up next to the desk and pointing their finger at the desk. A ray from their finger hits the table, where a text appears on top, saying ”select: table”. Three similar lines are extended towards the floor, wall, and ceiling. In pictures (b) and (c), the user is sitting on the desk, where a grid is seen covering the top surface of the desk. In the picture (b), the user is creating a cube object with a poking gesture with their right hand on the menu on the inside of their left wrist. In picture (c), the user is holding a real-world finger skateboard on their left hand while applying a pinch gesture with a right hand onto a virtual finger skateboard obstacle. In picture (d), the first-person view of the user is seen where a virtual printer twin is on top of the desk and the user is dragging the designed model with their left hand on the printer twin. The user is poking the button next to the digital printer twin labeled ”Start print”. In picture (e), the 3D-printed result can be seen on top of the desk.
@@ 1. Introduction
The introduction of Tinkercad AR [^1] https://www.tinkercad.com/ipad-app, Accessed: 16.07.2024 marked a notable progression by incorporating AR to view designs in physical spaces, albeit mainly for visualization. Applications such as Gravity Sketch [^2] https://www.gravitysketch.com/, Accessed: 16.07.2024 have begun to leverage virtual reality (VR) headsets and controllers to enable more immersive 3D sketching experiences. In recent HCI research, several notable advancements have emerged to enhance spatial and intuitive design interfaces. pARam (Stemasov et al., 2024) suggests that in-situ parametric design and customization can be effectively achieved through hand interactions and environmental considerations in AR. Additionally, BrickStARt (Stemasov et al., 2023a) introduces a novel approach by integrating tangible construction blocks with mixed reality (MR) to enable users to design and test physical artifacts directly in their intended locations. ProtoAR (Nebeling et al., 2018) addresses the challenges of AR content creation by combining physical prototyping with mobile authoring tools for rapid development and interaction. Furthermore, MixFab (Weichel et al., 2014) illustrates how MR can streamline personal fabrication by facilitating direct interaction with virtual objects within an AR environment. Collectively, these advancements highlight a trend towards more spatially aware and intuitive design interfaces, which are particularly beneficial for complex tasks that integrate digital and physical spaces.
@@
Despite these advancements, a significant gap remains in harnessing AR for interactive design that directly integrates with the user’s physical environment, balances expressivity and intuitiveness, supports seamless transitions from 3D modeling to 3D printing, and delivers all these capabilities in one cohesive package. Focusing on this gap, we introduce TinkerXR, an innovative AR interface that enables in-situ CAD and 3D printing for novices. Our system bridges the gap between simple, novice-friendly tools and immersive, spatial capabilities of advanced AR technologies, offering advanced spatial interaction techniques for designing 3D models directly in the physical environment using only hand controls. TinkerXR employs intuitive, Constructive Solid Geometry modeling with primitives similar to those in Tinkercad, but enhances the design experience by integrating these elements within the user’s physical environment through AR, making the design process more intuitive and adaptable to physical constraints and possibilities, while seamlessly integrating a 3D printing workflow.
@@
Our system utilizes a state-of-the-art headset (Meta Quest 3333https://www.meta.com/quest/quest-3/, Accessed: 16.07.2024) and its hand-gesture controls as well as passthrough features, enabling users to interact with and create 3D models directly atop their actual desks or on other flat surfaces like walls or the ground (Figure 3a). This setup not only facilitates intuitive interaction with the digital content but also allows users to accurately gauge the size and scale of their designs in relation to their immediate surroundings, enabling them to design in a way that considers and adapts to the space they’re working within.
Unlike other tools, TinkerXR is designed to recognize and adapt to real-world scene elements like furniture, enhancing its utility and applicability in everyday spaces, letting users turn any surface into their CAD workspace. Additionally, by designing around these setups, users can leverage features like snapping to grid and referencing smaller real-world objects, allowing users to design around them, making it easier to align and create models with precision, further enhancing the design process within their physical environment. Such features are critical in creating a truly immersive and intuitive design experience that seamlessly blends digital with physical.
It also incorporates depth occlusion, a technique that ensures digital objects are visually blocked by real-world objects based on their physical location and perspective, creating a more realistic and immersive experience. This is a notable advancement over other AR or VR systems, which may not achieve the same level of integration between the digital and physical worlds.
@@
Furthermore, TinkerXR extends the concept of AR in CAD beyond digital output to include actual manufacturing processes. It integrates seamlessly with 3D printing, incorporating a drag-and-drop metaphor that allows users to intuitively prepare models for printing (Figure 3c-e). Acting as a built-in slicing interface in AR, TinkerXR prepares models for printing while providing real-time feedback through a virtual 3D printer twin, allowing users to make adjustments and visualize the output. Users can also initiate the 3D printing operation directly from the system, ensuring a seamless transition from digital design to physical creation.
To validate the effectiveness of TinkerXR, we conducted a user study (n=10) with three design tasks. In addition to TinkerXR, we compared the results with Tinkercad as ground truth, using the same three tasks to assess performance across both systems. The study also included a survey to evaluate usability, user engagement, and the learning curve. Our findings indicate that TinkerXR enhances user engagement and supports novice users in design tasks. The results suggest high usability, with users reporting ease of use and a low learning curve.
@@
Overall, we present a novel design and fabrication system that integrates spatial and in-situ features with intuitive hand control schemes and a straightforward design logic tailored for novice users. By bridging the gap between the digital and physical worlds, our system enables seamless design-to-manufacturing workflows. Our contributions can be summarized as follows:
@@
  • We present TinkerXR, an innovative in-situ CAD system that enables hand-based CSG modeling directly in any physical environment, tailored for novices. The system’s novel contributions include interaction methods specifically designed for CSG design in real-world settings, such as snap-to-grid, designing on top of reference objects, and intuitive controls for applying traditional CAD functionalities with hand interactions.
@@
  • We integrate 3D printing workflows directly into the AR design process, allowing users to seamlessly transition from digital design to physical creation. The system includes a built-in slicer, real-time feedback via a virtual 3D printer twin, and the ability to initiate printing directly from the interface, offering a cohesive design-to-manufacturing pipeline.
  • We conduct a user study evaluating the effectiveness of TinkerXR when compared to Tinkercad, demonstrating high user engagement and immersiveness for novice users with no to little prior CAD experience.
@@
  • We make TinkerXR open-source [^4] The link was removed for anonymity., contributing to the research community by providing the system and codebase for further development, experimentation, and possible integration with additional fabrication technologies.
@@
Our approach simplifies geometric modeling in AR and supports a flexible methodology for designing in any place, smoothly transitioning from digital designs to physical outputs. By enhancing 3D creativity and accessibility through a context-aware immersive platform, we hope to advance the field of design and fabrication through AR.
@@ 2. Related Work
Situated Modeling (Lau et al., 2012) MixFab (Weichel et al., 2014) RoMA (Peng et al., 2018) DesignAR (Reipschläger and Dachselt, 2019) pARam (Stemasov et al., 2024) WindowShaping (Huo et al., 2017) TinkerXR
@@
Table 1. A comparison of features of AR-based CAD interfaces in the literature.
@@ 2.1. XR in Digital Design and Fabrication
VR tools have been mostly developed for digital design compared to fabrication. Tools like Gravity Sketch allow users to create 3D models in an immersive environment, providing a spatial understanding that traditional screen-based interfaces cannot offer. Systems such as VRSketchIn (Drey et al., 2020), AdaptiBrush (Rosales et al., 2021), and HandPainter (Jiang et al., 2021) serve as interim sketching tools that can facilitate the design workflow. Researchers also studied the effectiveness of CAD tools in VR (Bourdot et al., 2010; Feeman et al., 2018; Cordeiro et al., 2019). While these tools provide immersive spatial understanding for 3D modeling, they are typically limited to virtual-only environments, separating the design process from the physical context of deployment.
@@
In contrast to VR, AR operates directly in physical environment, allowing users to see their designs in their intended real-world context (Iyer et al., 2025; Dogangun et al., 2024; Dogan et al., 2023c). While head-mounted displays (HMDs) can be used with hand gestures (as seen in TinkerXR) or controllers in AR (Guo and Prabhakaran, 2022; Stemasov et al., 2024; Peng et al., 2018; Jasche and Ludwig, 2020), most of the previous AR systems rely on tablets with pen or finger interactions. For instance, massless [^5] https://massless.dev/, Accessed: 19.01.2025, focuses on reducing material waste during prototyping by overlaying design information directly onto the artifact’s deployment area, enhancing precision and minimizing errors due to physical constraints. Sketched Reality (Kaimoto et al., 2022) introduces an AR sketching tool that interacts with physically actuated robots, exploring the physics of designs and mechanisms. RealitySketch (Suzuki et al., 2020) introduces a dynamic and responsive interface for sketching interactive graphics and visualizations.
TinkerXR contributes to the vision of works such as (Wang et al., 2023; Alghofaili et al., 2023; Leiva et al., 2021, 2020), which streamline digital design and fabrication workflows. However, unlike these systems, TinkerXR uniquely integrates a 3D printing workflow directly within its AR interface, providing an all-in-one solution that eliminates the need for external slicing tools or connector software while supporting precise and fully-functional designs.
@@ 2.3. Novice Support Tools and Systems
Designing systems for novices requires consideration of simplicity, intuitiveness, and ease of learning. In the context of digital design and fabrication, this often translates into reducing the complexity of traditional CAD tools and making them more accessible. Tinkercad, for instance, has been successful in this regard by providing an entry-level CAD tool that embraces CSG modeling with simple geometric shapes and a drag-and-drop interface that is easy for beginners to grasp  (Stemasov et al., 2021). Other systems for design and fabrication using simplified core building blocks include FlatFitFab (McCrae et al., 2014), which connects planar parts for a 3D model, and similarly kyub (Baudisch et al., 2019), which uses simpler block-based design. PARTs (Hofmann et al., 2018) addresses the challenges faced by non-experts in 3D modeling by introducing a framework for expressing and reusing design intent. CraftML (Yeh and Kim, 2018) uses declarative syntax where the code dictates the output structure. Lee et al. (Lee et al., 2018, 2016) use pose and movement detection for personalized furniture design, and Tian et al. simplify analog fabrication with woodcutter  (Tian et al., 2018) and lathe (Tian et al., 2019) tools. Suzuki et al. (Suzuki et al., 2022) use virtual twins to visualize simulated behaviors when integrated with spatial references. Pfeuffer et al. (Pfeuffer et al., 2017) shows the use of virtual buttons on the wrists to access interactable menus.
Generative design has also been instrumental in achieving desired 3D outcomes with minimal effort. Some tools allow users to design and fabricate 3D models based on 2D sketches (Saul et al., 2010; Li et al., 2020; Kazi et al., 2017; Johnson et al., 2012; Igarashi et al., 1999), thereby reducing the complexity associated with directly creating models in 3D space. Other ways of generative design utilize speech inputs (Ballagas et al., 2019) and constrained parametric design (Shugrina et al., 2015). Furthermore, retrieving existing designs (Stemasov et al., 2020; Liang et al., 2022) and remixing (Follmer and Ishii, 2012; Follmer et al., 2010; Stemasov et al., 2023b; Roumen et al., 2018) enhance accessibility while balancing expressivity (Stemasov et al., 2021). ShapeFindAR (Stemasov et al., 2022) enables searching for future physical artifacts through spatial queries and text.

### v4

@@ Abstract.
Despite the growing accessibility of augmented reality (AR) for visualization, existing computer-aided design (CAD) systems remain confined to traditional screens or require complex setups or predefined parameters, limiting immersion and accessibility for novices. We present TinkerXR, an open-source AR interface enabling in-situ design and fabrication through Constructive Solid Geometry (CSG) modeling. TinkerXR operates solely with a headset and 3D printer, allowing users to design directly in and for their physical environments. By leveraging spatial awareness, depth occlusion, recognition of physical constraints, reference objects, and hand movement controls, TinkerXR enhances realism, precision, and ease of use. Its AR-based workflow integrates design and 3D printing with a drag-and-drop interface for printers’ virtual twins.
A user study comparing TinkerXR with Tinkercad shows that TinkerXR offers novices higher accessibility, engagement, and ease of use. Participants highlighted how designing directly in physical space made the process more intuitive. By bridging the gap between digital creation and physical output, TinkerXR aims to transform everyday spaces into expressive creative studios. We release TinkerXR as open source [^1] TinkerXR is available at http://tinkerxr.github.io. to encourage further exploration of accessible, spatially grounded CAD tools.
@@
journalyear: 2025copyright: ccconference: ACM Symposium on Computational Fabrication; November 20–21, 2025; Cambridge, MA, USAbooktitle: ACM Symposium on Computational Fabrication (SCF ’25), November 20–21, 2025, Cambridge, MA, USAdoi: 10.1145/3745778.3766651isbn: 979-8-4007-2034-5/2025/11ccs: Human-centered computing Interactive systems and tools
Figure 1. TinkerXR is an in-situ CAD and 3D printing AR interface. The figure illustrates key features such as spatial awareness, depth occlusion, and the use of physical objects and constraints for design. With Constructive Solid Geometry modeling, users can convert surfaces into interactive design canvases, and after designing, interact with 3D printers’ digital twins to prepare for and initiate fabrication. Examples of designs showcased in this figure include a pen holder, phone stand, cable organizer, plant pot holder, bottle mount, and key hook.:
@@ 1. Introduction
The introduction of Tinkercad AR [^2] https://www.tinkercad.com/ipad-app, Accessed: 16.07.2024 marked a notable progression by incorporating AR to view designs in physical spaces, albeit mainly for visualization. Applications such as Gravity Sketch [^3] https://www.gravitysketch.com/, Accessed: 16.07.2024 have begun to leverage virtual reality (VR) headsets and controllers to enable more immersive 3D sketching experiences. In recent HCI research, several notable advancements have emerged to enhance spatial design interfaces. pARam (Stemasov et al., 2024) suggests that in-situ parametric design and customization can be effectively achieved through hand interactions and environmental considerations in AR. Additionally, BrickStARt (Stemasov et al., 2023a) introduces a novel approach by integrating tangible construction blocks with mixed reality (MR) to enable users to design and test physical artifacts directly in their intended locations. ProtoAR (Nebeling et al., 2018) addresses the challenges of AR content creation by combining physical prototyping with mobile authoring tools for rapid development and interaction. Furthermore, MixFab (Weichel et al., 2014) illustrates how MR can streamline personal fabrication by facilitating direct interaction with virtual objects within an AR environment. Collectively, these advancements highlight a trend towards more spatially aware design interfaces, which are particularly beneficial for complex tasks that integrate digital and physical spaces (Suzuki et al., 2025).
@@
Despite these advancements, a significant gap remains in harnessing AR for interactive design that directly integrates with the user’s physical environment and supports seamless transitions from modeling in AR to 3D printing. Focusing on this gap, we introduce TinkerXR, an AR interface that enables in-situ CAD and 3D printing for novices. Our system bridges the gap between simple, novice-friendly tools and immersive, spatial capabilities of advanced AR technologies, offering spatial interaction techniques for designing 3D models directly in the physical environment using only hand controls. TinkerXR employs Constructive Solid Geometry modeling with primitives similar to those in Tinkercad, but enhances the design experience by integrating these elements within the user’s physical environment, integrating a 3D printing workflow, and making the design process more adaptable to the real world through AR.
@@
Our system utilizes a state-of-the-art headset [^4] Meta Quest 3: https://www.meta.com/quest/quest-3/, Accessed: 16.07.2025 and hand-gesture controls as well as passthrough features, enabling users to interact with and create 3D models directly atop their actual desks or on other flat surfaces like walls or the ground (Figure 3a). This setup facilitates interaction with the digital content, and also allows users to accurately gauge the size and scale of their designs in relation to their immediate surroundings.
Building upon prior work in environmentally-aware AR interfaces (Stemasov et al., 2024, 2023a; Weichel et al., 2014), TinkerXR is designed to leverage scene understanding to turn everyday surfaces into functional CAD workspaces. By incorporating spatial features such as snapping to environmental grids and referencing real-world objects, our system supports a spatially-aligned modeling workflow. To create a more coherent and visually realistic MR experience, it uses depth-sensing to ensure virtual objects are accurately occluded by real-world elements.
@@
Furthermore, TinkerXR extends the concept of AR in CAD beyond digital output to include actual manufacturing processes. It integrates directly with 3D printing by incorporating a drag-and-drop metaphor that allows users to quickly prepare models for fabrication. Acting as a built-in slicing interface in AR, TinkerXR prepares models for 3D printing while providing real-time feedback through a virtual 3D printer twin, which allows users to make adjustments and visualize the output in real-time.
To evaluate TinkerXR, we conducted a user study (n=10) comparing it against Tinkercad as a baseline for comparison. The participants completed three design tasks and a 3D print preparation task in both systems. We found no statistically significant difference in design completion time, while our integrated 3D printing workflow made the print preparation process faster than the traditional export-and-slice method. Subjectively, participants rated TinkerXR as having higher usability (SUS) (Brooke, 1996) and reported a lower mental workload (NASA-TLX) (Hart and Staveland, 1988), with the primary trade-off being a higher physical workload, which we discuss in the results.
@@
Overall, we present a novel design and fabrication system that integrates in-situ features with hand control schemes and a novice-friendly design logic. By bridging the digital and physical worlds, our system aims to enable seamless design-to-manufacturing workflows. We summarize our contributions as follows:
@@
  • We present TinkerXR, an in-situ CAD system that enables hand-based CSG modeling directly in any physical environment, tailored for novices. The system’s novel contributions include interaction methods specifically designed for CSG design in real-world settings, such as snap-to-grid, designing on top of reference objects, as well as novice-friendly controls for applying traditional CAD functionalities with hand interactions.
@@
  • We integrate 3D printing workflows directly into the AR design process, allowing users to seamlessly transition from digital design to physical creation. The system includes a built-in slicer, real-time feedback via a virtual 3D printer twin, and the ability to initiate printing directly from the immersive interface.
  • We present the results of a comparative user study that characterizes the trade-offs between TinkerXR and Tinkercad for novice users. Our findings show that TinkerXR achieves comparable design efficiency, streamlines the fabrication workflow, and reduces mental workload, at the cost of increased physical demand.
@@
  • We make TinkerXR open source [^5] TinkerXR is available at http://tinkerxr.github.io., contributing to the research community by providing the system and codebase for further development, experimentation, and possible integration with further fabrication technologies.
@@
Our approach simplifies AR-based geometric modeling and supports a flexible methodology for designing in any place, smoothly transitioning from digital designs to physical outputs. By enhancing 3D creativity through a context-aware immersive platform, we hope to advance the field of design and fabrication through AR.
@@ 2. Related Work
Table 1. Comparison of features of AR-based CAD interfaces reported in the literature.
Situated Modeling MixFab RoMA DesignAR pARam WindowShaping TinkerXR
(Lau et al., 2012) (Weichel et al., 2014) (Peng et al., 2018) (Reipschläger and Dachselt, 2019) (Stemasov et al., 2024) (Huo et al., 2017) [our system]
@@ 2.1. XR in Digital Design and Fabrication
VR tools have been mostly developed for digital design compared to fabrication. Tools like Gravity Sketch allow users to create 3D models in an immersive environment, providing a spatial understanding that traditional screen-based interfaces cannot offer. Systems such as VRSketchIn (Drey et al., 2020), AdaptiBrush (Rosales et al., 2021), HandPainter (Jiang et al., 2021), and HPIPainting (Cai et al., 2025) serve as interim sketching tools that can facilitate the design workflow. Researchers also studied the effectiveness of CAD tools in VR (Bourdot et al., 2010; Feeman et al., 2018; Cordeiro et al., 2019). While these tools provide immersive spatial understanding for 3D modeling, they are typically limited to virtual-only environments, separating the design process from the physical context of deployment.
@@
In contrast to VR, AR operates directly in physical environment, allowing users to see their designs in their intended real-world context (Iyer et al., 2025; Dogangun et al., 2025; Dogan et al., 2023c). While head-mounted displays (HMDs) can be used with hand gestures (as seen in TinkerXR) or controllers in AR (Guo and Prabhakaran, 2022; Stemasov et al., 2024; Peng et al., 2018; Jasche and Ludwig, 2020), most of the previous AR systems rely on tablets with pen or finger interactions. For instance, massless [^6] https://massless.dev/, Accessed: 19.01.2025, focuses on reducing material waste during prototyping by overlaying design information directly onto the artifact’s deployment area, enhancing precision and minimizing errors due to physical constraints. Sketched Reality (Kaimoto et al., 2022) introduces an AR sketching tool that interacts with physically actuated robots, exploring the physics of designs and mechanisms. RealitySketch (Suzuki et al., 2020) introduces a dynamic and responsive interface for sketching interactive graphics and visualizations.
TinkerXR contributes to the vision of works such as (Wang et al., 2023; Alghofaili et al., 2023; Leiva et al., 2021, 2020), which streamline digital design and fabrication workflows. Some systems focus on specialized fabrication processes, such as Palette-PrintAR (Lipkowitz and Desimone, 2024), which uses AR to design and simulate for multicolor resin 3D printing. Unlike these systems, TinkerXR uniquely integrates a 3D printing workflow directly within its AR interface, providing an all-in-one solution that eliminates the need for external slicing tools or connector software while supporting precise and fully-functional designs.
@@ 2.3. Novice Support Tools and Systems
Designing systems for novices requires consideration of simplicity and ease of learning. In the context of digital design and fabrication, this often translates into reducing the complexity of traditional CAD tools and making them more accessible. Tinkercad, for instance, has been successful in this regard by providing an entry-level CAD tool that embraces CSG modeling with simple geometric shapes and a drag-and-drop interface that is easy for beginners to grasp  (Stemasov et al., 2021). Other systems for design and fabrication using simplified core building blocks include FlatFitFab (McCrae et al., 2014), which connects planar parts for a 3D model, and similarly kyub (Baudisch et al., 2019), which uses simpler block-based design. PARTs (Hofmann et al., 2018) addresses the challenges faced by non-experts in 3D modeling by introducing a framework for expressing and reusing design intent. CraftML (Yeh and Kim, 2018) uses declarative syntax where the code dictates the output structure. Lee et al. (Lee et al., 2018, 2016) use pose and movement detection for personalized furniture design, and Tian et al. simplify analog fabrication with woodcutter  (Tian et al., 2018) and lathe (Tian et al., 2019) tools. In a similar vein, Draw2Cut (Gui et al., 2025) lowers the barrier to CNC milling by allowing novices to design by sketching directly on the physical material. Suzuki et al. (Suzuki et al., 2022) use virtual twins to visualize simulated behaviors when integrated with spatial references. Pfeuffer et al. (Pfeuffer et al., 2017) shows the use of virtual buttons on the wrists to access interactable menus.
Figure 2. Real-life workspace awareness. (a) Three cases for selecting the workspace in a scanned room. (b) Grid appearing on the selected workspace plane for the three corresponding cases. (c) Selecting the workspace based on a reference of a physical object. (d) Physical and digital objects rendered together with depth occlusion off (upper) versus depth occlusion on (lower).:
Generative design has also been instrumental in achieving desired 3D outcomes with minimal effort. Some tools allow users to design and fabricate 3D models based on 2D sketches (Saul et al., 2010; Li et al., 2020; Kazi et al., 2017; Johnson et al., 2012; Igarashi et al., 1999), thus reducing the complexity associated with directly creating models in 3D space. Other ways of generative design utilize speech inputs (Ballagas et al., 2019) and constrained parametric design (Shugrina et al., 2015). Further, retrieving existing designs (Stemasov et al., 2020; Liang et al., 2022) and remixing (Follmer and Ishii, 2012; Follmer et al., 2010; Stemasov et al., 2023b; Roumen et al., 2018) enhance accessibility while balancing expressivity (Stemasov et al., 2021). ShapeFindAR (Stemasov et al., 2022) enables searching for physical artifacts through spatial queries and text.